OpenStudy (anonymous):

Simplify the expression in exponential notation. a*a*a would this be 3a? Simplify the expression in exponential notation. 2*a*b*a*b 2ab^2? I haven't done these inawhile

OpenStudy (jdoe0001):

\(\bf \color{red}{a+a+a\implies 3a}\qquad \qquad a\times a\times a\implies a^3\\ \quad \\ 2\times a\times b\times a\times b\implies 2\times a\times a\times b\times b\implies 2a^2b^2\)

OpenStudy (anonymous):

thanks I thought so,but was unsure of my self